# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

### v2

@@ 3.1 Hierarchical Rubric Generation
Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository. The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels. Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation. The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details, with weights reflecting the relative importance of each component for understanding the repository
@@ Appendix A Example Generated Documentation
Appendix B Background and Related Work
B.1 Code Summarization and Documentation
B.1.1 Existing Approaches
Advances in natural language processing (NLP) have enabled significant progress in the task of code summarization. Existing methods [13, 21, 18, 16, 1, 4, 17, 29] primarily focus on function-level documentation, and can broadly be divided into two categories based on the type of context they exploit (see Table 2 for a comprehensive comparison):
Local-Level Context
Early neural approaches established the foundation for automated documentation. CodeBERT [9] demonstrated effective bimodal training on code and natural language, while more recent LLM applications have shown substantial improvements. For instance, [13] reported 20.6 BLEU scores across six languages using GPT-3, an 11.2% improvement over prior methods. DocuMint [21] further explored the use of small language models for Python docstring generation with promising results. However, these approaches primarily rely on local code context and do not leverage repository-level information to improve documentation quality.
Codebase-Level Context
Recognizing the limitations of function-only context, recent work has explored broader codebase information to enrich documentation. [18] demonstrated that incorporating class- and repository-level context can substantially improve summary quality through few-shot learning and retrieval-augmented generation. ProConSuL [16] incorporated project-level information via compiler analysis and call graphs, employing a two-phase training procedure to reduce hallucinations. RepoAgent [17] introduced comprehensive LLM-powered repository documentation, though it was limited in content diversity and architectural depth. DocAgent [29] advanced the field through multi-agent collaboration with topological processing, employing specialized agents that respect code dependencies to improve completeness and accuracy.
$\hookrightarrow$ Limitations and Our Motivation: Despite this progress, most approaches still target isolated snippets (e.g., functions or classes) and thus fail to capture the underlying architecture of the repository, as highlighted in Table 2. For a developer onboarding a completely new codebase, the primary challenge is not understanding what each individual function or class does, but rather grasping the high-level overview of the system, its dataflow and how different components interact. This underscores the need for an automatic methodology capable of generating documentation that spans from low-level details to high-level repository-wide perspectives, covering critical features and providing visual representations that support developer comprehension. Our approach, CodeWiki, addresses these limitations by generating true repository-level documentation while leveraging full codebase context and supporting multiple programming languages.
B.1.2 Existing Evaluation Methods
Most existing works evaluate the quality of generated documentation using textual similarity metrics such as BLEU or ROUGE. These metrics measure the overlap between generated text and reference ground-truth documentation in terms of lexical or token-level similarity.
$\hookrightarrow$ Limitations and Our Motivation: While these methods provide a lightweight quantitative measure, they often fail to capture deeper semantic aspects of the documentation, such as clarity, correctness, and completeness in explaining the underlying code behavior. Recent studies [23, 8] have shown that traditional metrics like BLEU and ROUGE do not consistently correlate with human judgment of documentation quality. This gap highlights the need for an assessment system that can fairly and accurately evaluate documentation, ensuring that the generated content truly supports developers in comprehending the system.
Table 2: Comparison of different code summarization approaches. Repo-Level Doc. indicates whether the approach generates repository-level documentation output. Codebase Context specifies whether the codebase information is used as input to generate the documentation.
Approach Repo-Level Doc. Codebase Context Multilingual DocuMint ASAP ProConSul DocAgent RepoAgent CodeWiki
B.2 Rubric-based evaluation
A repository can be described through multiple representations, each using different structures and presentation styles while sharing the same underlying architecture. This diversity makes it difficult to evaluate open-ended outputs automatically, as no single metric can reliably evaluate which version is better. Although human evaluation provides trustworthy results, it is resource-intensive and slow. To overcome these limitations and facilitate rapid assessment, many benchmarks [26, 14, 24, 2] often employ large language models (LLMs) as automated evaluators, using task-specific rubrics to judge the quality of generated outputs. In our CodeWikiBench benchmark, we similarly construct a set of hierarchical rubrics based on the official documentation, providing a structured and fine-grained framework to evaluate the quality and completeness of generated repository-level documentation.
@@ Appendix D Complementary Experimental Results
Repository Language LOC System Score (%) Coverage Improvement (%)

### v3

@@ 1 Introduction
2 Related Work
2.1 Code Intelligence and LLMs
The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks. Early work by von der Mosel et al. [47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3]. These established early transformer-based approaches for code understanding and generation.
Comprehensive surveys by Niu et al. [34] and Hou et al. [19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks. However, recent studies highlight important limitations: Li et al. [22] found that models often make confident decisions without sufficient reasoning, while Zheng et al. [57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52]. More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
2.2 Automated Documentation Generation
Documentation generation has emerged as a critical LLM application in software engineering. Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21]. While Sun et al. [45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context. Luo et al. [27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing. DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation. Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches. Roy et al. [41] and Mastropaolo et al. [29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25]. Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
2.3 Multi-Agent Systems for Software Engineering
Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving. AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks. MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology. In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle. The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification. However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
@@ 4.1 Hierarchical Rubric Generation
Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository. The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels. Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation. The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details.
@@ Appendix A Example Generated Documentation
Appendix B Related Work
B.1 Code Intelligence and LLMs
The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks. Early work by von der Mosel et al. [47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3]. These established early transformer-based approaches for code understanding and generation.
Comprehensive surveys by Niu et al. [34] and Hou et al. [19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks. However, recent studies highlight important limitations: Li et al. [22] found that models often make confident decisions without sufficient reasoning, while Zheng et al. [57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52]. More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
B.2 Automated Documentation Generation
Documentation generation has emerged as a critical LLM application in software engineering. Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21]. While Sun et al. [45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context. Luo et al. [27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing. DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation. Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches. Roy et al. [41] and Mastropaolo et al. [29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25]. Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
B.3 Multi-Agent Systems for Software Engineering
Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving. AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks. MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology. In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle. The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification. However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
@@ Appendix D Complementary Experimental Results
Repository Lang. LOC System Score (%) Cov. Imp. (%)